OpenStudy (anonymous):

while bicycling you increase your velocity from 30 km/h to 35 km/h. In the next second, you go from 35 km/h to 40 km/h. What is your acceleration?

OpenStudy (anonymous):

35km/h-30 km/h=5km/h=v time=1 second acceleration=v/t 5 x 3600/1000=18 18 m/s^2 is the acceleration

OpenStudy (anonymous):

first time (1 s) dv = 35 -30 = 5 km/h second time (2s) dv = 40 -35 = 5 km/h put the data into: acceleration = dv/time